Circular No. 7574

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001O IN ANONYMOUS GALAXY B. J. Boyle, S. M. Croom, I. J. Lewis, and J. Pogson, Anglo-Australian

Observatory; and K. Glazebrook, Johns Hopkins University, report the discovery of a supernova in an optical spectrum taken on Jan. 23.69 UT with the Anglo-Australian Telescope (+ 2^{o}-field fiber-fed multi-object spectrograph) in the course of the 2dF QSO Redshift Survey. The host galaxy, whose center is located at α = 10^{h}33^{m}36^{s}.36, δ = +0^{o}12′01′′.2 (equinox 2000.0), was selected for observation as being blue (U–B_{j} = −0.4, B_{j} = 20.3) and compact (as measured from 1984 U.K. Schmidt Telescope IIIa-J plates taken with GG395 and UG1 filters). SN 2001O is estimated to be located within 1′′ of the galaxy’s center (as the diameter of the 2dF optical fibers is 2′′). A magnitude of B_{j} ∼ 18.7 is derived for SN 2001O from the spectrum (a 1-hr integration at a low resolution of 0.7 nm). The spectrum of the host galaxy shows narrow emission lines in [O II], [O III], and the hydrogen Balmer series, indicating a redshift of z = 0.087 and suggesting the presence of a narrow-emission-line galaxy. Superimposed is the spectrum of a type-I supernova (probably type Ia or Ic) showing strong Si II and Ca II features at ∼ 610 and 390 nm (rest frame), respectively, suggesting that the supernova is before or at its peak luminosity.

V445 PUPPIS S. Kimeswenger, A. Bacher, C. Lederle, W. Kausch, W. Kapferer, and

G. E. Grömer, University of Innsbruck, write: “A set of four spectra (range 400–900 nm, resolution 0.26 nm) obtained with the Innsbruck 0.60-m telescope on Jan. 15.92 UT showed a continuum with a maximum intensity at 520 nm (cf. IAUC 7556). The spectrum still is dominated by Fe II lines. The most prominent feature, sharp-peaked at 660.8 nm, is not likely Ca I 657.3-nm (possibly N II 661.0-nm?). O VI at 682.5 and 708.2 nm are prominent (as in RR Tel). Although the features are blended, identification of He I at 402.5, 414.3, 443.7, and 587.5 nm, and He II at 541.1, 597.7, and 595.2 nm, seem to be likely. The FWHM of all features is > 1500 km/s.”

Visual magnitude estimates, provided in part by S. Otero, Buenos Aires, Argentina: Jan. 1.026 UT, 9.0 (M. Reszelski, Szamotuly, Poland); 4.243, 9.2 (J. G. de S. Aguiar, Campinas, Brazil); 9.796, 9.7 (A. Pearce, Nedlands, W. Australia); 13.083, 9.0 (P. Schmeer, Bischmisheim, Germany); 16.361, 9.7 (J. Bedient, Honolulu, HI); 18.367, 9.9 (Bedient); 20.028, 9.7 (Otero); 22.083, 9.6 (J. Rodriguez Freitas, Montevideo, Uruguay); 24.067, 9.6 (Otero); 26.047, 9.6 (V. Ladino, Barquisimeto, Venezuela).

2001 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green